OpenStudy (anonymous):

Which quadratic equation has roots 1/2 and 1/3? a. 5x2 - 5x - 2 = 0 c. 6x2 + 5x - 1 = 0 b. 5x2 - 5x + 1 = 0 d. 6x2 - 5x + 1 = 0

OpenStudy (anonymous):

d

OpenStudy (anonymous):

(x-(1/2))(x-(1/3)) = x^2 - 5x/6 + 1/6 Multiply by 6 gives d.

OpenStudy (anonymous):

6(1/2)^2 - 5(1/2) +1 6(1/4) - 5/2 + 2/2 3/2- 3/2 = 0 & 6(1/3)^2 - 5(1/3) +1 6/9 - 5/3 + 3/3 2/3 - 2/3 = 0